### v2

@@ Abstract
With the advancements in quantum computing, transitioning to post-quantum cryptography (PQC) is becoming increasingly critical to maintain the security of modern systems. This paper introduces a formal definition of the cryptographic migration problem and explores its complexities using a suitable directed graph model. Characteristics of the resulting migration graphs are analyzed and trade-offs discussed. By using classical mathematical results from combinatorics, probability theory and combinatorial analysis, we assess the challenges of migrating “random” large cryptographic IT-infrastructures. We show that any sufficiently large migration project that follows our model has an intrinsic complexity, either due to many dependent (comparatively easy) migration steps or due to at least one complicated migration step. This proves that in a suitable sense cryptographic migration is hard in general. Furthermore, we analyze the proposed model with respect to practical applicability and explain the difficulties that emerge when we try to model real-world migration projects.
Keywords:
@@
Cryptographic migration, post-quantum cryptography, hardness results
1 Introduction
@@
Cryptography is fundamental ensuring information security in today’s digital world. Despite advances in cryptographic technologies, the process of migrating existing systems and IT-infrastructures to new standards presents significant challenges [5]. Examples include the transition from IPv4 to IPv6, which took more than 20 years due to the complexity, and technologies such as gateways and NAT became popular migration tools to circumvent the need to actually migrate from IPv4 to IPv6. A cryptographic example is the update of key sizes (RSA keys got bigger and so did the integration problems) or the transition between different hash functions (SHA-1 to the SHA-2 family to the SHA-3 family) with very slow adoption of the SHA-3 family as it is seen as an add-on or fallback rather than a replacement. For a rapid transition, it seems that there needs to be a strong incentive in terms of the benefits of adopting the new technology. One notable example of quicker migration, at least in the broad public, is the transition from the Data Encryption Standard (DES) to the Advanced Encryption Standard (AES), which was driven by the need for stronger security measures due to DES’s vulnerability to brute force attacks [22].
@@
This issue is particularly prominent in the current transition towards PQC, which aims to protect data against future quantum computing threats. The German Federal Office for Information Security (BSI) uses the work-hypothesis for the high-security sector that with significant probability there is a cryptographically relevant quantum computer in the early 2030s [11]. Thus, there are numerous practical post-quantum migrations going on world-wide. The reason is that we are already running-out of time: Sensitive (e. g. classified) data has long confidentiality requirements (e. g. 25 years). In addition migrating the IT-infrastructure correspondingly takes a considerable amount of time. This runs against estimates on how long it will take to build a cryptanalytically relevant quantum-computer, c. f. Mosca’s famous inequality [16]. To understand the structure of cryptographic migration and to potentially improve the practicability of real-world migrations, we will look at the cryptographic migration problem, i. e. the task of migrating a given IT-infrastructure from one cryptographic algorithm to another one, from a mathematical point of view. Despite many common-sense approaches being applied to tackle PQC migration, there remains a significant gap in scientifically grounded methodologies. We address that gap by providing a formal model that lays the foundation for a more rigorous understanding and systematic approach to the complexities of cryptographic migration.
@@
This paper is structured as follows: An overview of related work can be found in Section 2. Our first contribution is to provide a unified and consistent notation by introducing a formal model of cryptographic migration in Section 3. In Section 4 we prove that migration is generally hard in a suitable sense: Namely, we show that any “random”, reasonably large migration project has an inherent complexity. This is our second contribution. Sections 5 and 6 further describe our work in terms of its usage and its potential applications, which shall serve as a foundation for upcoming real-world migrations – our third contribution.
2 Background and Related Work
We will model the cryptographic migration problem as a suitable graph problem. Directed graphs have a long and rich history in computer sciences, see e. g. classical standard references such as [4, 14]. They are used in a wide range of domains due to their ability to model asymmetric relationships and dependencies. For instance, they are used to represent the information exchange in distributed systems [1], the dynamics of the world wide web [28], and communication protocols in network design [26]. To properly place directed graphs in the context of cryptographic migration, we first examine the current state: After the announcements of the first post-quantum public key algorithms for consideration as an international standard by the National Institute of Standards and Technology (NIST) [2], the transition to PQC has started in practice: in the US, there is a large project run by NIST with a large consortium of industry-leading companies that explores the transition to PQC [6]. In Europe, the BSI published a guideline concerning post-quantum migration [10] and also included certain post-quantum algorithms in their recommended suite of cryptographic algorithms [12]. The Dutch government also published a “PQC Migration Handbook” containing guidelines for migrating to PQC [3]. Despite these world-wide activities, we found little research on the structure of migration projects. In fact, many people consider the actual migration of soft- and hardware components somewhat as a best-practice task, and people legitimately ask “Where Is the Research on Cryptographic Transition and Agility?” as David Ott did at RWC2022 (see also [20]). But is migration really a best-practice task? Shouldn’t there be structural reasons why cryptographic migration is so hard to achieve in practice? There are a few informal models of the migration problem (e. g. [21, 18, 27, 9]) but formal treatment is, as far as the authors know, still pending. To objectify this knowledge, we conducted a systematic literature review on this subject [17] without finding a formal model comparable to the one proposed here. This does not completely rule out that there is literature around we did not consider, but we assume that our model can substantially contribute to the formal basics of cryptographic migration.
3 A Formal Model of the Migration Problem

### v3

@@ Abstract
With the advancement of quantum computing, the transition to post-quantum cryptography (PQC) is becoming increasingly critical to maintain the security of modern dependable infrastructural systems. This paper presents a novel approach to gain insight into the structure of cryptographic migration problems, using a semi-formal model to capture the inherent dependencies and complexities of such transitions. Using classical mathematical results from combinatorics, probability theory, and combinatorial analysis, we assess the challenges of migrating large cryptographic IT-infrastructures and prove that – in a suitable sense – cryptographic migration has a certain expected complexity. Furthermore, we analyze the proposed model in terms of real-world patterns as well as its practical applicability, and discuss difficulties that arise when trying to model real-world migration projects. This work sets the stage for future advances in both the theoretical understanding and practical implementation of cryptographic migration strategies in the post-quantum era.
Index Terms:
@@
Cryptographic migration, post-quantum cryptography, semi-formal model
I Introduction
@@
Cryptography is fundamental to ensuring the information security in today’s digital world and is thus one important cornerstone of many modern dependable systems. Despite advances in cryptographic technologies, the process of migrating existing systems and IT-infrastructures to new standards presents significant challenges [3]. One example is the transition from IPv4 to IPv6, which took more than 20 years and promoted technologies such as gateways and NAT to circumvent the actual migration instead. A cryptographic example is the update of key sizes (RSA keys got bigger and so did the integration problems) or the transition between different hash functions (SHA-1 to the SHA-2 family to the SHA-3 family) with slow adoption of the SHA-3 family as SHA-2 seems still sufficient for most tasks. It seems that a rapid transition requires a strong incentive in terms of the benefits of adopting the new technology. One notable example of quicker migration, at least in the broad public, is the transition from the Data Encryption Standard (DES) to the Advanced Encryption Standard (AES), which was driven by the need for stronger security measures due to DES’s vulnerability to brute force attacks [20].
@@
The issue of migration is particularly prominent in the current transition towards PQC, which aims to protect data against future quantum computing threats. The German Federal Office for Information Security (BSI) uses the work-hypothesis for the high-security sector that there likely exists a cryptographically relevant quantum computer in the early 2030s [8]. Thus, there are numerous practical post-quantum migrations going on world-wide. The reason is that we are already running-out of time: Sensitive (e. g. classified) data often has long confidentiality requirements (e. g. 25 years). In addition, migrating the IT-infrastructure correspondingly takes a considerable amount of time [24]. This runs against estimates on how long it will take to build a crypto analytically relevant quantum-computer, c. f. Mosca’s famous inequality [13].
To be able to conduct such migrations as efficiently as possible we aim to model the PQC migration problem mathematically and look for potential optimal strategies. As two relevant parts, i.e. the modeling of IT networks and dependencies, are often depicted as graphs [11], we also examine the migration problem in the context of a graph-based model. By this systematic analysis, we hope to better understand the structural problems and find strategies necessary for building tools to assist with conducting the actual migrations. Our findings show that usually the systems’ dependencies of a to-be-migrated infrastructure necessitate a compatible migration order, otherwise functionality is lost. Within this order there are always sets of systems which need to be migrated together – so called migration clusters. As soon as the systems under evaluation exhibit non-trivial complexity, we have many or large clusters, and these clusters are expected to be interdependent, making migration difficult. To facilitate the process and still allow for rapid migration, we see a need for tools that build on our results to find and visualise the optimal migration graphs.
@@
After an overview of related work in Section II, we present our first contribution in Section III. It consists of a formalization approach for cryptographic migration, a discussion of the architectural structure of the resulting model, and explores typical patterns that occur in practice. In Section IV we prove that any (reasonably large) migration project has an expected intrinsic complexity as our second contribution. Section V further describes our model in terms of its usage and its potential applications which shall serve as a foundation for tools which assist the upcoming real-world migrations – our third contribution.
II Background and Related Work
After the announcements of the first post-quantum public key algorithms for consideration as an international standard by the National Institute of Standards and Technology (NIST) [1], the transition to PQC has started in practice: in the US, there is a project run by NIST with a large consortium of industry-leading companies that explores the transition to PQC [3]. In Europe, the German BSI published a guideline concerning post-quantum migration [7] and also included certain post-quantum algorithms in their recommended suite of cryptographic algorithms [9]. The Dutch government also published a “PQC Migration Handbook” containing guidelines for migrating to PQC [2].
Despite these world-wide activities, we found little research on both, the general structure of migration projects and how to suitably model cryptographic migration. In fact, many people consider migrating of soft- and hardware components as a best-practice task. So, people legitimately ask “Where Is the Research on Cryptographic Transition and Agility?” as David Ott did at RWC2022 (see also [18]). Shouldn’t there be structural reasons why cryptographic migration is so hard to achieve in practice? We will show in Section IV that there is!
Recent efforts have focused on consolidating the state-of-the-art in cryptographic migration [14] and systematically organizing knowledge on cryptographic agility [16]. Furthermore, frameworks arise that aim to support the migration process as in [25], and the development of tools to assist in the migration process [11]. The latter follows – as we do – a graph-based approach, providing a solid starting point for PQC migration by introducing a practical framework focused on cryptographic and data inventories and dependency analysis within enterprise systems. However, the analysis in [11] primarily addresses operational aspects and could benefit from a more formal approach. Building on this groundwork, our efforts enhance the theoretical rigor by formalizing migration challenges through mathematical models. Generalizing the approach might even address broader migration scenarios beyond PQC-specific contexts.
Also to be mentioned are few informal models of the migration problem (e. g. [19, 15, 23, 6]). However, formal treatment is, as far as the authors know, still pending. We believe that our migration model together with its analysis can substantially contribute to the formal basics of cryptographic migration.
III A Model for the Cryptographic Migration Problem
Any IT-infrastructure we migrate consists of a number of migration components or simply components which need to be migrated. Such migration components can be software libraries, whole software packages, full operating systems, network protocols, etc. just to name a few. For a more thorough study of which components we might encounter in practice see Section III-B. The migration problem is now to migrate all migration components. However, not every component can be migrated at every point in time, because there might be dependencies between the components. For example, a given software package can only be migrated to post-quantum if its used cryptographic libraries can. Another example is the migration of a communicating software component, which can only be migrated if the corresponding counterparts of the communication are migrated at the same time. In this case the migration can be done (if there are no other dependencies) if all communicating parties of this single migration component are migrated at the same time. We call this as stated in the introduction in Section I a migration cluster. Components without dependencies outside its migration cluster are called migratable components. The migration process looks now as follows:
  1. 1.
    In every migration step we select, following a suitable migration strategy, a migratable component.
  2. 2.
    We then compute its migration cluster, i.e. possible interdependent components that need to be migrated at the same point in time.
  3. 3.
    Then, the actual migration of the migration cluster is performed and the successfully migrated migration cluster is removed from our model (together with all in-going dependencies).
  4. 4.
    This process is repeated until there is no migration component left and we have then solved the migration problem for this specific IT-infrastructure.
The whole model is now captured in a directed graph whose vertices are the migration components and the edges are the dependencies between components, the migration graph of the migration problem, see Appendix -A for the formal details.
Note that a migration graph’s structure will vary significantly depending on the concrete IT-infrastructure and the level of model we have in mind, c.f. Section III-A. But also the abstraction layer and the type of migration lead to many design choices of the model. The resulting graphs will show parallels to those that already exist in several disciplines of computer science. Those comprise flowcharts, call graphs, and finite-state machines in programming, network diagrams, or software dependency graphs, just to name a few. Additionally, current efforts in the creation of Software Bills of Materials (SBOMs) or Cryptography Bills of Materials (CBOMs) and their respective tooling add to that collection of suitable graphs.
We expect that post-quantum migration is the most significant type of cryptographic migration of IT-infrastructures in the near future [14]. Its specific properties introduce aspects that were less prevalent in the classical cryptography world. One example are hybrid cryptographic schemes. Depending on the specific model, a hybrid solution may just be seen as a single cryptographic primitive or cipher suite. Yet, modeling detailed software dependencies or taking dependencies on space/throughput requirements into account might have a significant effect.
Post-quantum migration also emphasizes the importance of cryptographic agility (see for example [18]). Many communication protocols used to be optimized for a limited set of somewhat similar algorithms. The introduction of crypto-agile mechanisms that support various kinds of cryptographic primitives and, sometimes arbitrary, combinations of these, will lead to an increase in complexity. In a graph which only models systems and their protocol-related communication relations in between, one may only treat this implicitly. Yet again, while graphs that take other dependencies into account usually become harder to model, they sometimes better reflect the complicated circumstances of the real world.
Also, regulatory requirements of post-quantum solutions and different views on current and future algorithm recommendations may become significantly more complicated, especially in a multinational context.
III-A Real-World Migration Graphs
Today’s large scale IT-infrastructures are both complex and highly diverse, and examining models to create a suitable graph from a real-world IT-infrastructure is an open research topic. The aggregation of the necessary data that is used to derive the graph is highly individual and specific to the given migration goal. We state a few concrete examples of such modeling attempts:
Network migration:
When talking about network migration it seems intuitive to think of the components to represent computers, servers, and similar devices. Much thought has to be taken into the dependencies. These could, e. g. model communication paths for specific (cryptographic) protocols. For the same network, all connections between machines may be interesting regardless of the specific protocol, illustrating all the relations between machines in the network. Depending on the objective one may study several graphs representing different kinds of communication relations each – for example individual graphs for protocols like TLS/HTTPS, SSH, or IPsec/IKEv2.
System migration:
Different graphs may be modeling a specific system, e. g. the migration components being a set of software-artifacts within a firmware or operating system. It is unclear at which granularity such modeling may be done. We expect that relations may be handled similarly to SBOMs, showing dependencies between software, libraries, and so on. This may even go into the details of a specific software project showing relations between different modules, library calls, etc.
Full-scale migration:
For a system administrator, a tool focusing on the high-level network view may be enough, while programmers may profit from call-graph-like models. A full view on all abstraction levels may be provided by a combination of different graphs, e. g. a component being linked to another graph depicting the same component in a different context. This leads us to having differing graph models of the same infrastructure, just depending on our perspective.
Note that when we consider classical, naturally grown enterprise networks, the graphs will probably be both: large and complex. Also, they will have many interdependent systems. It is beyond our current model to capture time dependent changes in the migration graph. But even when we look at a snap-shot of a migration graph, the design choices for modeling real-world IT-infrastructure graphs are vast and we discovered a variety of possible ways to model such IT-infrastructures.
III-B Real-World Patterns
$v$
$v_{1}$$v_{2}$$v_{3}$$w$
$v$$w_{2}$$w_{1}$$w_{3}$
$v$$w$
$v_{1}$$v_{2}$$v_{3}$
Figure 1: Dependency sub-graphs of real-world migration attempts: 1(a) Single component with no dependencies. 1(b) Many components depend on a single component. 1(c) One component depends on many components. 1(d) Mutual dependencies between two components. 1(e) Higher-order cycles.
API-compatible change of algorithms:
Operating System dependencies:
Certain components $v_{1}$, $v_{2}$, $v_{3}$, $\dots$, for example software or libraries, can only be migrated when the underlying execution platform $w$ supports the newly migrated component. This often occurs when one version of an application is updated but the new version only runs on a newer version of the operating system. This corresponds to case Fig. 1(b). In the context of post-quantum migration, this could be the replacement of a necessary cryptographic hardware resource running a certain operating system with a PQC-ready version of it.
Regulatory requirements:
Communication protocols:
When migrating communication protocols from one version to another, it is vital that all communicating parties in the migration cluster are migrated at the same time. This is in particular a problem when there are many parties involved and only one part of the own IT-infrastructure may be updated conveniently, e. g. by software configuration management tools with external dependencies. This problem often leads to backwards compatibility as a feature, but likely also leads to the possibility of downgrading attacks. This corresponds to case Fig. 1(d). In the context of post-quantum migration this could be the replacement of TLS by a post-quantum-ready version of it.
Dependency on space/throughput requirements:
Occasionally, there are requirements on the amount of memory, the capabilities of computing platforms, or network throughput restrictions. This corresponds to cases Figs. 1(d) and 1(e). In the context of post-quantum migration this applies when for example the key sizes of post-quantum algorithms are much larger than those of classical algorithms.
Dependencies of software-packages within distributions:
When considering software-packages and their dependencies, the graph formed this way is a directed acyclic graph with no mutual dependencies. This graph represents combinations of cases illustrated in Figs. 1(a)1(b), and 1(c). The migration of such a case can be done corresponding to the canonical topological order of this graph. In the context of post-quantum migration this could be the replacement of cryptographic libraries with post-quantum-ready versions.
Cyclically depending assets:
The draft of ETSI’s “Repeatable Framework for Quantum-safe Migrations” [6] explains that if for example three cryptographic assets are in a dependency relationship with each other, functionality issues may occur when a single asset is migrated. The functionality issues might cascade and affect all assets in the circle, even the migrated one. This corresponds to case Fig. 1(e) but the problem is mitigated in our model as we assume that such circular clusters of assets can only be migrated when all assets are migrated at the same time. In the context of post-quantum migration this might happen for example when we observe circularly cross-signed certificates. Such circular cases are rare in practice, because they create dependencies that are avoided in software architecture and engineering. Also, developers can address those kinds of dependencies by introducing adapter nodes in one direction which allow to break the dependency cycle.
IV Assessing the Difficulty of the Migration Problem
We now show that, expectedly, any sufficiently large migration project has an intrinsic complexity. Specifically we show:
Theorem 1.
Expectedly we have for migration projects with $n$ migration components within our model:
  1. 1.
    The number of migration steps $M$ is expectedly $\mathsf{E}(M)\sim n/\mathop{\text{log}}\nolimits n$ with standard deviation $\mathsf{SD}(M)\sim\sqrt{n}/\mathop{\text{log}}\nolimits n$ and the expected size of each migration cluster is $\mathop{\text{log}}\nolimits n$.
  2. 2.
    $\frac{n}{\mathop{\text{log}}\nolimits n}+\frac{\sqrt{2n\mathop{\text{log}}\nolimits s}}{\mathop{\text{log}}\nolimits n}.$
  3. 3.
    If the set of migratable components is at each step bound by a smooth function ${f\colon\mathbb{R}\rightarrow\mathbb{R}}$ in a real variable with $f\in o(n)$, then we expect $n/f(n)$ non-independent migration steps that successively depend on each other.
Proof.
Since every migration project decomposes disjointly into its migration clusters, we now study the expected properties of these migration clusters. Let $n$ be the number of components. There are $B_{n}$ different possible partitions of $n$ components in disjoint subsets, where $B_{n}$ is the $n$th Bell number. Denote by $M$ the number of disjoint subsets of a randomly chosen partition.
  1. 1.
    The statement on the expectation and the standard deviation of $M$ is a summary of [17] and Chapter 4, Theorem 1.1 in [22]. Furthermore, the expected maximum size of a migration cluster is, for growing $n$, asymptotic to $er-\mathop{\text{log}}\nolimits\sqrt{r}$, where $r\cdot\mathop{\text{exp}}\nolimits(r)=n$ and $r=\mathop{\text{log}}\nolimits n-(1+o(1))\mathop{\text{log}}\nolimits\mathop{\text{log}}\nolimits n\sim\mathop{\text{log}}\nolimits n$ [21]. Results on the speed of convergence are beyond the scope of this article.
  2. 2.
    $\leq\sum_{i=1}^{s}\mathsf{E}(\mathop{\text{exp}}\nolimits(tM_{i}))$
    $=s\mathop{\text{exp}}\nolimits(t^{2}\sigma^{2}/2)$
    Here, the last equality follows from the fact that the moment generating function $\mathsf{E}(\mathop{\text{exp}}\nolimits(tM_{i}))$ for a normal distribution with mean zero is equal to $\mathop{\text{exp}}\nolimits(t^{2}\sigma^{2}/2)$ and in particular independent of the index-variable $i$. Rewriting this, $\mathsf{E}(X)\leq\frac{\mathop{\text{log}}\nolimits s}{t}+\frac{t\sigma^{2}}{2}$. This bound is minimal at $\frac{\sqrt{2\mathop{\text{log}}\nolimits s}}{\sigma}$ since its derivative in direction $t$ is $\frac{\sigma^{2}}{2}-\frac{\mathop{\text{log}}\nolimits s}{t^{2}}$. Thus, $\mathsf{E}(X)\leq\sigma\sqrt{2\mathop{\text{log}}\nolimits s}$ and the result follows.
  3. 3.
    We study the dependencies between the migration clusters. According to the first statement of this Theorem, we consider the directed acyclic graph resulting from the migration graph on $n$ components modulo its strongly connected components with expectedly $n^{\prime}=n/\mathop{\text{log}}\nolimits n$ vertices. Since we are considering a single migration project, we can assume that this graph is connected. We now use the standard Erdős-Rényi model [5], where this graph is constructed by connecting the components independently at random with a fixed probability $p$. One can ensure that the result is acyclic by defining any linear order on the vertices and include a directed edge only if it obeys the selected ordering. Each such edge is included in the graph with probability $p_{n}$, independently from every other edge. For such graphs, we have by Theorem 4.1 in [10], a variant of Theorem 1 in [4] that for such graphs, for growing $n$, the probability that $G$ is connected is given by
    $\begin{cases}0&\text{, if $c_{n}\to-\infty$}\\ e^{-e^{-c}}&\text{, if $c_{n}\to c$ constant}\\ 1&\text{, if $c_{n}\to\infty$}\\ \end{cases}$ (1)
    where $p_{n}=\frac{\mathop{\text{log}}\nolimits n+c_{n}}{n}$. Note that we can directly apply this Theorem to our directed acyclic graph situation, since such a graph is connected if and only if the undirected variant of it is connected.
    As our goal is to prove a lower bound on the number of edges in the condensation graph (with $n^{\prime}$ vertices!) in our model. We assume now the sparsest case $c_{n^{\prime}}\to c$ constant for which the graph is connected with high probability. Note that the doubly-exponential term in Equation (1) approaches $1$ already for comparatively small $c$ very quickly. For example, already for $c=10$, the value of $e^{-e^{-c}}\geq 0.99995$. But also for any $c>0$, we obtain for the number of edges in the condensation graph expectedly asymptotically
    $\frac{\mathop{\text{log}}\nolimits n^{\prime}+c_{n^{\prime}}}{n^{\prime}}\cdot\frac{n^{\prime}\cdot(n^{\prime}-1)}{2}\sim\frac{1}{2}n^{\prime}\mathop{\text{log}}\nolimits n^{\prime}.$
    That means that even in the sparsest case, we expect a considerable number of dependencies between the migration clusters. Furthermore, if we bound the number of components with a common ancestor of distance $k$, i. e. #$\#(v\to^{k}V)$ by a small function $f(n)=o(n)$, e. g. $f(n)=\mathop{\text{log}}\nolimits n$, we obtain by Theorem 4 in Appendix -D for the condensation graph expectedly $d(G)\geq\frac{n}{f(n)\mathop{\text{log}}\nolimits n}$, that is, large chains of dependencies of the migration clusters.
This completes the proof of our main Theorem, proving that in a suitable sense we expect for any large migration project an expected intrinsic complexity with many small interdependent clusters. ∎
We thus showed that informally we have
Fact.
Any sufficiently large migration project has an intrinsic complexity due to many dependent (comparatively small) migration clusters. The migration of these clusters either takes many successive steps or includes at least one particularly difficult one.
[…]
TABLE I: Numerical evaluation of the expected behavior of migration clusters for migration projects with $n$ components rounded to the nearest integer. We see that one does not expect large mutual dependencies but the expected number of independent clusters is vast and the chain-dependencies are also large when we bound the number of migration clusters with fixed canonical migration time by $f(n)=\mathop{\text{log}}\nolimits n$ or $f(n)=\sqrt{n}/\mathop{\text{log}}\nolimits n$, respectively.
We conclude that even though the model does not represent the exact structure of every migration project, it gives a first sound basis for understanding the structure of migration projects in general. We summarize these results with a numerical evaluation of the relevant estimates in Table I with notation from Appendix -A.
V Practical Implications
Ott et al. define the term cryptographic agility as “the ability to make a transition from one cryptographic algorithm or implementation to another” [18]. Adopting cryptographic agility in that way ensures the flexibility to transition between algorithms or implementations in response to evolving threats. Thus, in order to implement cryptographic agility we need to prepare cryptographic migration. In turn, for preparing migration, we need a step-by-step migration plan. This plan must allow careful re-evaluation of dependencies and system integrity at each stage. This kind of approach is in line with making security solutions sustainable by allowing to adapt to changing threats and technologies. We have shown that the discussed migration approach based on graph models allows step-by-step evaluation and reevaluation throughout the whole process.
Furthermore, we have shown – as stated in Theorem 1 – big migration projects are intrinsically complex. These projects require many successive steps and in-depth strategic planning. Migration graphs can help in understanding this complexity involved and in considering the time needed to migrate, but also specify the number of components and the various dependencies. As we have seen, these dependencies can exist between single components, between interdependent clusters of components, and between external factors such as regulatory requirements, legal issues, or third-party components. The complexity of migrations requires expert knowledge and sufficient dedicated resources, resulting in significant costs. Economic implications and the risks of both, migrating and not migrating to PQC, must be evaluated. It is also crucial for stakeholders, policy makers, and non-technical decision makers to grasp this complexity and to develop appropriate organizational measures to address it. The model we propose can serve as a starting point to identify the highest priority parts of the system, e. g., by identifying the parts which are at the beginning of a migration chain, or which have the highest impact. We expect that from there, an iterative approach works best and we intend to investigate this in future work.
Given this general structure of practical migration, we think that our model can significantly support and enhance existing migration strategies and tools, particularly in the standardization of migration projects through guidelines, rules, and best practices. These models can optimize the process, provide practical examples for guidance, and reduce cost and effort. Initial steps towards such standardization include frameworks like CARAF, a risk management approach for post-quantum migration [12]. CARAF incorporates a threat and a risk assessment stage where our migration graphs can prove valuable. During threat assessment, migration clusters identified using our approach can be used to identify vulnerable subsystems. If one of the systems in the cluster is vulnerable, so is the cluster in its entirety. Likewise, during risk assessment, information about which systems need to be migrated together is helpful to establish cost and effort estimates. Consequently, we believe our approach complements and strengthens existing methods in the field, such as those explored by Näther et al. [14].
VI Conclusion
In this work, we used classical mathematical results from graph theory, probability theory, and combinatorial analysis to assess the difficulty of a very practical problem: Cryptographic migration of large IT-infrastructures. We formulated a formal model of cryptographic migratability in terms of directed graphs and showed several expected properties of these graphs. This approach not only elevates the discussion to a scientific level but also establishes a foundation for further research on more efficient migration strategies. In summary, we showed that migration projects always have an intrinsic complexity due to many dependent (comparatively small) migration clusters which (successively) depend on each other.
Still, there is a lot of work to be done, when it comes to practical migration projects: It is not clear how to obtain useful migration graphs in practice, such that they are neither trivial nor too vast. Analyzing how to model real-world IT-infrastructures or relevant subsets will clearly be a challenging task. Also, there is clearly room for improvements: As migration problems typically come with some kind of measurable cost per migration step, a refinement of our model would result in migration graphs where optimal migration strategies are not as easy to obtain as here. One further important open question is also the development of migration graphs and the changes in such graphs over time. Our work sets the stage for future research aimed at refining these models, integrating them with real-world data, and ultimately creating more adaptive, user-friendly tools that support real-world migration. We therefore aim here to scientifically substantiate migration approaches. By emphasizing the relevance of this formal approach, we hope to support the development of more robust migration strategies and (open-source) tooling that can keep pace with the evolving landscape of cryptographic standards. Future work should apply our model to real-world migration projects. The hope is to get a better insight of the complexity such projects not just in general, as discussed here, but for specific practically occurring cases. This might help to better understand real-world cryptographic migration processes, specifically in the context of PQC migration.
Acknowledgments
The authors thank Tobias J. Bauer for his valuable comments during the genesis of this article.
-A Migration Graphs